# Characterizing Multi-Hunk Patches: Divergence, Proximity, and LLM Repair Challenges

### v1

@@ I Introduction
Automated Program Repair (APR) has received significant attention from the research community [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. APR techniques have progressed from heuristic-based methods [19, 20, 21, 22] to deep learning approaches that learn repair patterns from large codebases [15, 16, 17, 10, 18], achieving promising results on single-hunk bugs [16, 17, 18, 10, 5, 4, 23, 24, 25, 26, 27]. Despite recent progress, most APR research remains focused on single-hunk scenarios [16, 24, 17, 18, 10], overlooking multi-hunk bugs, defects that require edits across disjoint code regions. These are both common in practice [28] and harder to fix due to their semantic interdependencies. While the problem has been acknowledged [28], only a few studies have directly addressed it [29, 28, 30, 31, 32, 33], and they have demonstrated limited effectiveness.
@@
Recent advancements in large language models (LLMs) have shown promise in transforming several software engineering tasks, including program repair [34]. Trained on massive datasets comprising code and natural language, LLMs such as GPT-4 and LLaMA have demonstrated the ability to understand and generate code [35, 36, 37, 34]. While these models have already been employed to address single-hunk bugs with notable success, their potential to handle multi-hunk code repair remains underexplored.
@@
Early studies on LLM-based multi-hunk repair include tools such as AlphaRepair [35], RepairAgent [38], and fine-tuning strategies [39, 40]. Although these approaches demonstrate that LLMs can extend beyond single-hunk scenarios, the challenges of reasoning about semantically interdependent edits distributed across the codebase remain unaddressed.
@@
We identify four challenges limiting progress. First, dataset variation impedes consistency: prior studies evaluate on synthetic benchmarks such as LMDefects [41], algorithmically curated tasks (EvalRepair-C++/Java) [40], and real-world datasets such as Defects4J [42], often under varying assumptions, input formats, and test harnesses. Second, bug-type ambiguity obscures evaluation: multi-hunk bugs span a wide spectrum, from edits in non-contiguous regions within a single function to interdependent changes across multiple files. Yet prior work does not systematically characterize this heterogeneity, often conflating distinct forms of multi-location repair such as multi-line, multi-function, and multi-file edits. Third, reporting inconsistencies hinder meaningful attribution: existing techniques often report aggregate metrics that combine single- and multi-hunk results [40, 38], making it difficult to isolate and evaluate their effectiveness specifically on multi-hunk repairs. Fourth, existing evaluations lack patch complexity-aware metrics that quantify differences in hunks within a patch.
@@ II Characterization of Multi-hunk Patches
Although multi-hunk patches have been studied [28, 33, 38], their structural and spatial properties and how they affect automated repair remain poorly understood.
@@
$h_{i}=(\text{loc}_{i},\text{content}_{i},\text{file}_{i},\text{method}_{i})$
@@ II-A Hunk Divergence
To compute lexical distance ($D_{\text{lex}}$), we tokenize the diff exactly as it appears in the patch, preserving the ordering of added and deleted lines. For hunks involving both additions and deletions, we concatenate the deleted lines (before the change) and the added lines (after the change) prior to tokenization. In cases where a hunk contains only additions or only deletions, the token sequence consists solely of the respective lines as shown in the diff. We use BLEU [43] to quantify lexical similarity between token sequences of code fragments, following prior work [44, 45, 46, 47, 48, 17, 10, 49, 50].
@@
[…]
@@
$D_{\text{file}}(h_{i},h_{j})=\begin{cases}0&\text{if }f_{i}=f_{j}\\ 1-\dfrac{|\texttt{LCP}(f_{i},f_{j})|}{\max\big{(}|f_{i}|,\ |f_{j}|\big{)}}&\text{otherwise}\end{cases}$
Here, $\texttt{LCP}(f_{i},f_{j})$ denotes the length of the longest common prefix between the directory sequences $f_{i}$ and $f_{j}$, the two distinct buggy files, and ||$\cdot$|| ⋅ | denotes the number of directory components in a given path. This formulation produces a normalized estimate of file-level distance: file paths that share a deeper hierarchical structure yield lower divergence values. This provides a continuous, repository hierarchy-aware quantification of inter-file separation that is both bounded and independent of path depth. It captures whether a pair of hunks is confined to the same file or spans across distinct files, serving as a proxy for the inter-file coordination effort required during multi-hunk repair.
@@ 27      $lcpLen\leftarrow\textsc{LCP}(f_{1},\ f_{2})$
[…]
@@ II-B Spatial Proximity
$\mathrm{SP}(P)$ $=\exists\mathcal{P}\;\text{such that}\;\forall h_{i}\in P,\;\mathcal{P}_{i}=\mathcal{P}$
$\mathrm{LCP}(P)$ $=\min_{i\neq j}\left|\text{LCP}(\mathcal{P}_{i},\mathcal{P}_{j})\right|$
@@
$\mathcal{C}(P)=\begin{cases}\texttt{Nucleus}&\text{if }\mathrm{SF}(P)\land\mathrm{SM}(P)\\ \texttt{Cluster}&\text{if }\mathrm{SF}(P)\land\lnot\mathrm{SM}(P)\\ \texttt{Orbit}&\text{if }\lnot\mathrm{SF}(P)\land\mathrm{SP}(P)\\ \texttt{Sprawl}&\text{if }\lnot\mathrm{SF}(P)\land\lnot\mathrm{SP}(P)\land\mathrm{LCP}(P)>\lambda\\ \texttt{Fragment}&\text{if }\lnot\mathrm{SF}(P)\land\lnot\mathrm{SP}(P)\land\mathrm{LCP}(P)\leq\lambda\\ \end{cases}$
@@ II-C Hunk4J: A Dataset of Multi-Hunk Bugs
To enable principled evaluations, we introduce Hunk4J, a curated dataset of multi-hunk bugs derived from Defects4J [51]. Defects4J includes reproducible, test-suite backed bugs from open-source Java projects across diverse domains. We analyze all the 835 bugs in Defects4J [^1] We use Defects4J version 2.0.1 for our analysis. and identify those whose developer-written patches contain multi-hunk fixes (Definition 2). We found that 372 out of 835 bugs (44.6%) are multi-hunk. These 372 multi-hunk bugs form the basis of our Hunk4J for further analysis. Beyond extracting patch-level characteristics, we augment Hunk4J with natural language context for each bug. Specifically, we collect bug reports, titles, and summaries from official project repositories and issue trackers, including Apache JIRA, GitHub Issues, and SourceForge. This enriched metadata facilitates prompt construction and retrieval-augmented strategies for LLM-based program repair.
@@ II-D Hunk Characterization
We categorize multi-hunk bugs along two axes: the number of change regions (hunks) and the number of files touched in Table I. Single-file multi-hunk bugs dominate, with 244 cases where changes remain confined within a single file—this includes 140 two-hunks, 55 with three hunks, and 49 with four or more. These reflect intra-file repair scenarios, requiring local reasoning. In contrast, 128 bugs span multiple files, representing more fragmented, distributed repairs. Notably, 68 of these involve four or more hunks, indicating considerable dispersion across the codebase. For instance, multi-hunk patch $P$ for Jsoup_56 [52] touches 10 hunks across 5 files.
To complement our earlier categorization, Table II presents project-specific statistics capturing how multi-hunk bugs manifest across different codebases. In 15 out of 17 projects, the median number of hunks per bug is at most 3, suggesting that the majority of bugs involve relatively localized changes. However, outliers exist, Jsoup_87 [53] has 47 hunks. In contrast, Chart_4 [54] involves a multi-hunk fix with only two hunks confined to a single method, which is simpler to reason about. Across all 17 projects, the median number of files modified per bug does not exceed 2; nonetheless, notable exceptions are present. For instance, JacksonDatabind_103 [55] spans 26 hunks across 16 files, modifying exception handling across the codebase. Conversely, certain projects such as Lang, Collections, and JacksonXml contain no multi-file bugs.
@@
Table III categorizes multi-hunk patches based on their spatial proximity, revealing a distinct skew toward more localized fixes. The Cluster class, comprising 185 bugs, is the most prevalent and includes patches where multiple hunks occur within the same file but in different methods. This is followed by the Orbit class (67 bugs), which captures cases where hunks span multiple files within the same package, and the Nucleus class (59 bugs), where all hunks are confined to a single method. Mockito_6 [56] from the Cluster class contains 20 hunks, all confined to a single file yet distributed across multiple methods, demonstrating file-level cohesion despite method-level dispersion. In contrast, the Sprawl (50 bugs) and Fragment (11 bugs) classes represent increasingly dispersed patches. The Fragment category consists of patches where hunks are scattered across unrelated packages. An example from Fragment class is Closure_144 [57], in which the hunks are scattered across unrelated packages, highlighting the low spatial locality. The average hunk divergence increases monotonically with spatial dispersion. Fragment patches, though rare, exhibit the highest divergence (0.7372), underscoring their complexity.
@@ III-A Benchmarking Framework: Birch
Implementation. Birch is implemented in Python and supports both proprietary models and open-source models via Ollama [58]. We use litellm [59] to interface with LLM APIs across providers. Prompts are templated using TOML, enabling integration of new prompt variants. We employ JavaParser to parse ASTs, which are used to compute node-level distances between hunks, derive the tree diameter, and extract enclosing scopes such as methods and classes. For sparse retrieval, we use BM25; for dense retrieval, we employ text-embedding-text-small-3 or Mini-LM, storing embeddings in a FAISS [60] index for nearest‐neighbor lookup.
@@ III-B Models
This study conducts a comparative evaluation of six models, selected to represent both proprietary and open-source variants with diverse architectural characteristics. The proprietary models include Gemini-2.5-Flash (1M token context window) from Google, evaluated in non-reasoning mode via the official Gemini API, and Nova-Pro (300K tokens) from Amazon. From OpenAI, we evaluate both a reasoning-capable model, o4-mini (200K tokens), and a non-reasoning variant, GPT-4.1 (1M tokens), accessed directly via the OpenAI API. Parameter counts for proprietary models are not publicly disclosed. The open-source models comprise mistral-large-2407 (128K tokens, 123B parameters) from Mistral AI and LLaMA3.3 (128K tokens, 70B parameters) from Meta. All models, except those from OpenAI and Google, were evaluated using full-parameter inference endpoints through Amazon Bedrock [61]. Our goal is not to perform exhaustive model benchmarking. Rather, we aim to assess the generalizability of our findings across LLMs with varying architectures.
@@ III-C Evaluation Metrics
We assess patch correctness based on whether the generated code parses, compiles, and passes all tests provided by the benchmark suite. Following prior work [62], we report Plausible@1, defined as the fraction of bugs for which the top-1 generated plausible patch passes the complete test suite:
[…]
@@
We generate one patch per bug using standard decoding settings (e.g., temperature = 0.0) for all models except o4-mini, whose API has a fixed temperature of 1.0. While decoding is configured to be deterministic at the API level, generation remains non-deterministic in practice due to low-level system factors such as thread scheduling, memory layout, and backend randomness [63].
@@ III-D Effectiveness of LLMs on Multi-Hunk Repair (RQ1)
Overall, despite recent advances, LLMs still struggle with multi-hunk bug repair, as even the best model fails to fix over 70% of cases. The higher accuracy of o4-mini highlights the benefit of using reasoning-capable models with greater test-time compute compared to non-reasoning models. However, the limited repair accuracy even in the best-performing model suggests that increasing test-time compute alone may not be sufficient, particularly for complex repair tasks [64].

### v2

@@ I Introduction
Automated Program Repair (APR) has received significant attention from the research community [yasunaga:drrepair:icml20, cure-program-repair-icse-2021, codebert-java-sstubs-21, recoder-fse-2021, rewardrepair-icse-2022, hata2019learning, watson:lit-review-dl4se:tosem:22, le:srevey-dl4se:csur20, tufano:tosem:19, reptory, coconut, sequencer, zhang:apr-survey:tosem23, huang:apr-survey:acm24, deepdelta, hoppity, glance, katana]. APR techniques have progressed from heuristic-based methods [legoues:apr:acm19, legoues:genprog:tse11, weimer:everaging-program-equivalence:ase13, kim:par-template-based-apr:icse13] to deep learning approaches that learn repair patterns from large codebases [deepdelta, hoppity, glance, reptory, katana], achieving promising results on single-hunk bugs [hoppity, glance, katana, reptory, rewardrepair-icse-2022, recoder-fse-2021, li:dlfix:icse20, xuan:nopol:tse16, jiang:simfix:issta18, zhu:tare:icse23, yang:transplantfix:ase22]. Despite recent progress, most APR research remains focused on single-hunk scenarios [hoppity, xuan:nopol:tse16, glance, katana, reptory], overlooking multi-hunk bugs, defects that require edits across disjoint code regions. These are both common in practice [hercules] and harder to fix due to their semantic interdependencies. While the problem has been acknowledged [hercules], only a few studies have directly addressed it [mechtaev:angelix-multi-hunk:icse16, hercules, yuan:arjae:19, wong:varfix:fse21, dear, iter], and they have demonstrated limited effectiveness.
@@
Recent advancements in large language models (LLMs) have shown promise in transforming several software engineering tasks, including program repair [cedar]. Trained on massive datasets comprising code and natural language, LLMs such as GPT-4 and LLaMA have demonstrated the ability to understand and generate code [xia:alpharepair:fse22, hossain:llm-bug-localization-and-repair:fse24, zhang:critical-review-chatgpt-and-apr:arxiv23, cedar]. While these models have already been employed to address single-hunk bugs with notable success, their potential to handle multi-hunk code repair remains underexplored.
@@
Early studies on LLM-based multi-hunk repair include tools such as AlphaRepair [xia:alpharepair:fse22], RepairAgent [repair-agent], and fine-tuning strategies [huang:empirical-fine-tuning-for-apr:ase23, yang:fine-tuning-for-repair-with-llm:arxiv24]. Although these approaches demonstrate that LLMs can extend beyond single-hunk scenarios, the challenges of reasoning about semantically interdependent edits distributed across the codebase remain unaddressed.
@@
We identify four challenges limiting progress. First, dataset variation impedes consistency: prior studies evaluate on synthetic benchmarks such as LMDefects [fan:code-repair-with-llm:icse23], algorithmically curated tasks (EvalRepair-C++/Java) [yang:fine-tuning-for-repair-with-llm:arxiv24], and real-world datasets such as Defects4J [srepair:arxiv24], often under varying assumptions, input formats, and test harnesses. Second, bug-type ambiguity obscures evaluation: multi-hunk bugs span a wide spectrum, from edits in non-contiguous regions within a single function to interdependent changes across multiple files. Yet prior work does not systematically characterize this heterogeneity, often conflating distinct forms of multi-location repair such as multi-line, multi-function, and multi-file edits. Third, reporting inconsistencies hinder meaningful attribution: existing techniques often report aggregate metrics that combine single- and multi-hunk results [yang:fine-tuning-for-repair-with-llm:arxiv24, repair-agent], making it difficult to isolate and evaluate their effectiveness specifically on multi-hunk repairs. Fourth, existing evaluations lack patch complexity-aware metrics that quantify differences in hunks within a patch.
@@ II Characterization of Multi-hunk Patches
Although multi-hunk patches have been studied [hercules, iter, repair-agent], their structural and spatial properties and how they affect automated repair remain poorly understood.
@@
$h_{i}=(\text{loc}_{i},\text{content}_{i},\text{file}_{i},\text{method}_{i},\text{pkg}_{i})$
@@
  • $\text{pkg}_{i}$: package path as a sequence of directory segments.
@@ II-A Hunk Divergence
To compute lexical distance ($D_{\text{lex}}$), we tokenize the diff exactly as it appears in the patch, preserving the ordering of added and deleted lines. For hunks involving both additions and deletions, we concatenate the deleted lines (before the change) and the added lines (after the change) prior to tokenization. In cases where a hunk contains only additions or only deletions, the token sequence consists solely of the respective lines as shown in the diff. We use BLEU [bleu] to quantify lexical similarity between token sequences of code fragments, following prior work [zhou:vulnerability-repair:icse24, shirafuji:code-repair:icast23, gu:deep-api-learning:fse16, jiang:commit-message-nmt:ase17, watson:assert-statement:icse20, glance, reptory, apr-survey:tosem24, koutcheme:distance-measure-repair:icer23].
@@
$D_{\text{ast}}=\begin{cases}\frac{\ln(1+\textsc{ASTDist}(n_{1},n_{2}))}{\ln(1+\textsc{TreeDiameter}(\mathcal{T}_{f}))}&\text{if }\textsc{TreeDiameter}(\mathcal{T}_{f})>0\\ 0&\text{otherwise}\end{cases}$
@@
[…]
Here, $\texttt{LCP}(f_{i},f_{j})$ denotes the length of the longest common prefix between the directory sequences $f_{i}$ and $f_{j}$, the two distinct buggy files, and $|\cdot|$ denotes the number of directory components in a given path. This formulation produces a normalized estimate of file-level distance: file paths that share a deeper hierarchical structure yield lower divergence values. This provides a continuous, repository hierarchy-aware quantification of inter-file separation that is both bounded and independent of path depth. It captures whether a pair of hunks is confined to the same file or spans across distinct files, serving as a proxy for the inter-file coordination effort required during multi-hunk repair.
@@ 27 $lcpLen\leftarrow\textsc{LCP}(f_{1},\ f_{2})$
In Algorithm 1, we show the steps for computing the pairwise divergence $D(h_{1},h_{2})$ between two hunks. Line 1 computes lexical divergence using BLEU score. Lines 11 implement the conditional logic for determining whether the hunks are co-located in the same file. Line 1 handles the intra-file case: the structural divergence is computed as the normalized distance between AST nodes within the same file, using the AST’s diameter for normalization. File-level divergence is omitted by setting $D_{\text{file}}=0$. In contrast, the cross-file case (beginning at Line 1) assigns $D_{\text{ast}}=1$.0 to denote maximal syntactic divergence across disjoint ASTs. File-level divergence is then computed using the normalized longest common prefix (LCP) between the two file paths. The scaling factor $\gamma$ is set to $2.0$ to amplify the influence of file-level separation. This dynamic adjustment of $\gamma$ reflects the increased effort required when developers must coordinate edits across multiple files, ensuring that file-level fragmentation contributes more significantly to the overall divergence score in multi-hunk patches.
@@ II-B Spatial Proximity
$\mathrm{SP}(P)$ $=\exists p\;\text{such that}\;\forall h_{i}\in P,\;\text{pkg}_{i}=p$
$\mathrm{LCP}(P)$ $=\min_{i\neq j}\left|\text{LCP}(\text{pkg}_{i},\text{pkg}_{j})\right|$
@@
[…]
@@ II-C Hunk4J: A Dataset of Multi-Hunk Bugs
To enable principled evaluations, we introduce Hunk4J, a curated dataset of multi-hunk bugs derived from Defects4J [defects4j]. Defects4J includes reproducible, test-suite backed bugs from open-source Java projects across diverse domains. We analyze all the 835 bugs in Defects4J [^1] We use Defects4J version 2.0.1 for our analysis. and identify those whose developer-written patches contain multi-hunk fixes (Definition 2). We found that 372 out of 835 bugs (44.6%) are multi-hunk. These 372 multi-hunk bugs form the basis of our Hunk4J for further analysis. Beyond extracting patch-level characteristics, we augment Hunk4J with natural language context for each bug. Specifically, we collect bug reports, titles, and summaries from official project repositories and issue trackers, including Apache JIRA, GitHub Issues, and SourceForge. This enriched metadata facilitates prompt construction and retrieval-augmented strategies for LLM-based program repair.
@@ II-D Hunk Characterization
We categorize multi-hunk bugs along two axes: the number of change regions (hunks) and the number of files touched in Table I. Single-file multi-hunk bugs dominate, with 244 cases where changes remain confined within a single file—this includes 140 two-hunks, 55 with three hunks, and 49 with four or more. These reflect intra-file repair scenarios, requiring local reasoning. In contrast, 128 bugs span multiple files, representing more fragmented, distributed repairs. Notably, 68 of these involve four or more hunks, indicating considerable dispersion across the codebase. For instance, multi-hunk patch $P$ for Jsoup_56 [jsoup56-multi-hunk-patch] touches 10 hunks across 5 files.
To complement our earlier categorization, Table II presents project-specific statistics capturing how multi-hunk bugs manifest across different codebases. In 15 out of 17 projects, the median number of hunks per bug is at most 3, suggesting that the majority of bugs involve relatively localized changes. However, outliers exist, Jsoup_87 [jsoup87-multi-hunk-patch] has 47 hunks. In contrast, Chart_4 [chart4-multi-hunk-patch] involves a multi-hunk fix with only two hunks confined to a single method, which is simpler to reason about. Across all 17 projects, the median number of files modified per bug does not exceed 2; nonetheless, notable exceptions are present. For instance, JacksonDatabind_103 [jackson103-multi-hunk-patch] spans 26 hunks across 16 files, modifying exception handling across the codebase. Conversely, certain projects such as Lang, Collections, and JacksonXml contain no multi-file bugs.
@@
Table III categorizes multi-hunk patches based on their spatial proximity, revealing a distinct skew toward more localized fixes. The Cluster class, comprising 185 bugs, is the most prevalent and includes patches where multiple hunks occur within the same file but in different methods. This is followed by the Orbit class (67 bugs), which captures cases where hunks span multiple files within the same package, and the Nucleus class (59 bugs), where all hunks are confined to a single method. Mockito_6 [mockito6-multi-hunk-patch] from the Cluster class contains 20 hunks, all confined to a single file yet distributed across multiple methods, demonstrating file-level cohesion despite method-level dispersion. In contrast, the Sprawl (50 bugs) and Fragment (11 bugs) classes represent increasingly dispersed patches. The Fragment category consists of patches where hunks are scattered across unrelated packages. An example from Fragment class is Closure_144 [closure144-multi-hunk-patch], in which the hunks are scattered across unrelated packages, highlighting the low spatial locality. The average hunk divergence increases monotonically with spatial dispersion. Fragment patches, though rare, exhibit the highest divergence (0.7372), underscoring their complexity.
@@ III-A Benchmarking Framework: Birch
Implementation. Birch is implemented in Python and supports both proprietary models and open-source models via Ollama [ollama]. We use litellm [litellm] to interface with LLM APIs across providers. Prompts are templated using TOML, enabling integration of new prompt variants. We employ JavaParser to parse ASTs, which are used to compute node-level distances between hunks, derive the tree diameter, and extract enclosing scopes such as methods and classes. For sparse retrieval, we use BM25; for dense retrieval, we employ text-embedding-text-small-3 or Mini-LM, storing embeddings in a FAISS [johnson:billion-scale-search:ieeetran19] index for nearest‐neighbor lookup.
@@ III-B Models
This study conducts a comparative evaluation of six models, selected to represent both proprietary and open-source variants with diverse architectural characteristics. The proprietary models include Gemini-2.5-Flash (1M token context window) from Google, evaluated in non-reasoning mode via the official Gemini API, and Nova-Pro (300K tokens) from Amazon. From OpenAI, we evaluate both a reasoning-capable model, o4-mini (200K tokens), and a non-reasoning variant, GPT-4.1 (1M tokens), accessed directly via the OpenAI API. Parameter counts for proprietary models are not publicly disclosed. The open-source models comprise mistral-large-2407 (128K tokens, 123B parameters) from Mistral AI and LLaMA3.3 (128K tokens, 70B parameters) from Meta. All models, except those from OpenAI and Google, were evaluated using full-parameter inference endpoints through Amazon Bedrock [amazon-bedrock]. Our goal is not to perform exhaustive model benchmarking. Rather, we aim to assess the generalizability of our findings across LLMs with varying architectures.
@@ III-C Evaluation Metrics
We assess patch correctness based on whether the generated code parses, compiles, and passes all tests provided by the benchmark suite. Following prior work [repairbench], we report Plausible@1, defined as the fraction of bugs for which the top-1 generated plausible patch passes the complete test suite:
$\text{Plausible@1}=\frac{1}{|\mathcal{B}|}\sum_{b\in\mathcal{B}}\mathrm{TestPass}(\mathsf{patch}_{1}^{b})$
@@
We generate one patch per bug using standard decoding settings (e.g., temperature = 0.0) for all models except o4-mini, whose API has a fixed temperature of 1.0. While decoding is configured to be deterministic at the API level, generation remains non-deterministic in practice due to low-level system factors such as thread scheduling, memory layout, and backend randomness [shuyin:non-determinism-of-chatgpt:tosem25].
@@ III-D Effectiveness of LLMs on Multi-Hunk Repair (RQ1)
Overall, despite recent advances, LLMs still struggle with multi-hunk bug repair, as even the best model fails to fix over 70% of cases. The higher accuracy of o4-mini highlights the benefit of using reasoning-capable models with greater test-time compute compared to non-reasoning models. However, the limited repair accuracy even in the best-performing model suggests that increasing test-time compute alone may not be sufficient, particularly for complex repair tasks [snell:scaling-llm-test-time:iclr25].